OpenStudy (anonymous):

What are the solutions of 2x2 + 8x = –26? The last one I need help understanding. I don't want the answer and I did most of the work I'm stuck at a part tho.

OpenStudy (anonymous):

So far i got \[\frac{-8 \pm \sqrt{8^2-4(2)(26)}} {2(2)}\]=\[\frac{-8 \pm \sqrt{64-208}} {4}\]=\[\frac{-8 \pm \sqrt{-144}} {4}\] what I suppose to do from here? Answer, A.\[-2\pm3i\]B.\[-2\pm i\]C.\[2\pm 6i\]D.\[-3\pm 2i\]

OpenStudy (anonymous):

Stuck here- \[\sqrt{-144} = \sqrt{-1\times144}\] i think it goes like this \[\sqrt{-144}= \sqrt{-1\times12\times12}=\sqrt{-144}= \sqrt{-1}\times \sqrt{12}\times \sqrt{12}\] I think this is right so far.. :/

OpenStudy (anonymous):

@jim_thompson5910 I know I've asked alot of questions today on this but can you help me on my last on please :/

OpenStudy (anonymous):

*one

jimthompson5910 (jim_thompson5910):

why is this question in the photography section?

jimthompson5910 (jim_thompson5910):

its fine and we can work here, just curious

OpenStudy (anonymous):

I dont know I feel stupid I clicked on the wrong section and just now saw that :(

jimthompson5910 (jim_thompson5910):

that's fine, we can stay here

jimthompson5910 (jim_thompson5910):

2x^2 + 8x = –26 2x^2 + 8x + 26 = 0 2(x^2 + 4x + 13) = 0 x^2 + 4x + 13 = 0

OpenStudy (anonymous):

oh so i did that all wrong?

jimthompson5910 (jim_thompson5910):

You can plug in a = 2, b = 8, c = 26 into the quadratic formula, but I find it easier to simplify as much as possible first

jimthompson5910 (jim_thompson5910):

let's go your way, but going either way gets you the same answer

jimthompson5910 (jim_thompson5910):

because it all reduces anyway

OpenStudy (anonymous):

oh okay

jimthompson5910 (jim_thompson5910):

2x^2 + 8x = –26 2x^2 + 8x + 26 = 0 a = 2 b = 8 c = 26 \[\Large x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[\Large x = \frac{-(8)\pm\sqrt{(8)^2-4(2)(26)}}{2(2)}\] \[\Large x = \frac{-8\pm\sqrt{64-208}}{4}\] \[\Large x = \frac{-8\pm\sqrt{-144}}{4}\] \[\Large x = \frac{-8+\sqrt{-144}}{4} \ \text{or} \ x = \frac{-8-\sqrt{-144}}{4}\] \[\Large x = \frac{-8+12i}{4} \ \text{or} \ x = \frac{-8-12i}{4}\] \[\Large x = -2+3i \ \text{or} \ x = -2-3i\]

jimthompson5910 (jim_thompson5910):

If we used a = 1, b = 4, and c = 13, then we would get the same answers, so either path works

OpenStudy (anonymous):

oh okay thanks again for all the help :/ I don't understand much of this math that well.

jimthompson5910 (jim_thompson5910):

just keep practicing and it will (hopefully) click better

OpenStudy (anonymous):

lol I'm sure it will, I'm normal A student in math but doing a grade ahead of my grade level on line is not so easy on your own lol.

jimthompson5910 (jim_thompson5910):

dang, that's impressive lol well keep up the good work

jimthompson5910 (jim_thompson5910):

and yes, it's tough on your own without a teacher to help

OpenStudy (anonymous):

thanks and yeah it is but I'm doing pretty good I finish my first semester in a totally time of 11weeks today.

jimthompson5910 (jim_thompson5910):

that's great, so it's like an accelerated course?

OpenStudy (anonymous):

I got held back in 1st grade so I'm a year behind. And so I decided half way through this year that I'd make that year up and graduate next year. I started a little late but I'm making it work. But yeah I'm in 10th and doing 11th grade classes at home online on top of 10th schooling.

jimthompson5910 (jim_thompson5910):

well I think a lot of people do since some kids aren't ready for school, but not many would take on two grades to catch up

OpenStudy (anonymous):

Yeah. Everyone thinks I'm crazy and I can't do it. But what the heck If it means I finish high school in 3 years and not 4. And I graduate with my age group then so be it. And it will look really good on collage applications lol.

jimthompson5910 (jim_thompson5910):

yeah for sure, it will show that you're a very hard worker, which is what I would look for most in a potential employee

jimthompson5910 (jim_thompson5910):

lol and that's fuel to show them wrong and that you can do it

OpenStudy (anonymous):

Yeah well nice chatting and thanks for helping me get to the answer and not just giving it to me. But I got to go finish these math essay questions I have to do now. then go take my semester 1 finals exam. And I can't not tell you how thankful I am! :D

jimthompson5910 (jim_thompson5910):

glad to be of help, good luck with the rest of it and your classes (and 2 grade levels)

OpenStudy (anonymous):

Thanks! :)

candycove (candycove):

o-o